Searches related to a body weighs 2000 lbf when exposed to the standard earth gravity g=32.174 ft/s2, what is the mass

1
Expert's answer
2021-02-05T04:54:47-0500

P=m*g

P= 2000 lbf = 8896.443 N

g= 32.174 ft/s2

=9.8m/s2


m=P/g

=8896.443/9.8

=907.8003 kg